solve using the elimination method.
-x-y=3
2x-y=-3

Answer by mananth(16946)

-x-y=3
2x-y=-3
subtract the equations
-x-y-(2x-y)=3-(-3)
-x-y-2x+y=6
-3x=6
x=6/-3
x=-2
-x-y=3
-(-2)-y=3
+2-y=3
-y=3-2
-y=1
y=-1